Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers.
510 and 92

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Expressing 510 and 92 as product of its prime factors, we get,
510 =2 × 3 × 17 × 5 × 1
92 = 2× 2 × 23 × 1
Therefore,LCM(510, 92) = 2 × 2 × 3 × 5 × 17 × 23 = 23460
And HCF ( 510,92) = 2
Verification
Now, product of 510 and 92 = 510× 92 = 46920
And product of HCF and LCM= 23460 × 2 = 46920
Hence, LCM × HCF = product of the 510 and 92.

​Thanks